If s = 2t3 − 21t2 +60t describes the displacement of a body at time t from a fixed point O on a straight line, find:


(A). s when t = 3 seconds


b)         the values of t when the body is at rest


c)         the initial velocity of the body


d)         an expression for the acceleration of the body at time t


e)         the initial acceleration of the body.   


1
Expert's answer
2021-04-15T07:40:08-0400

"s(t)=2t^3-21t^2+60t"


a) "s(3)=2\cdot 3^3-21\cdot 3^2+60\cdot 3=54-189+180=45"


b) The body is at rest when "v(t)=0" .

"v(t)=s^\prime (t)=6t^2-42t+60=0"

"6t^2-42t+60=0"

"t^2-7t+10=0"

"D=(-7)^2-4\cdot10=9"

"t_{1,2}=\frac{7\pm \sqrt{9}}{2}"

"t_1=2" and "t_2=5"


c) "v(0)=6\cdot 0^2-42\cdot 0+60=60"


d) "a(t)=s^{\prime \prime}(t)=v^\prime (t)=12t-42"


e) "a(0)=12\cdot 0-42=-42"


Answer: a) "s(3)=45" ; b) "t_1=2" and "t_2=5" ; c) "v(0)=60" ; d) "a(t)=12t-42" ; e) "a(0)=-42" .